OpenStudy (anonymous):

2x(x + 18) - x(4x - 2)

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

do u know how to do this

OpenStudy (anonymous):

do you mean 2x(x + 18) - x(4x - 2)=0

OpenStudy (stamp):

2x(x+18) = 2xx + 2x18 = 2xx + 36x x(4x-2) = 4xx - 2x 2xx + 36x - (4xx -2x) 2xx + 36x - 4xx -(-2x) 2xx + 36x - 4xx + 2x 2xx - 4xx + 36x + 2x

OpenStudy (anonymous):

it dose not have the zero at the end

OpenStudy (anonymous):

u have to Simplify it

OpenStudy (anonymous):

let me see..

OpenStudy (stamp):

2xx - 4xx + 36x + 2x -2xx + 38x\[-2x^2+38x=x(-2x+38)\]

OpenStudy (anonymous):

-2x2 + 38x -2x2 + 38 -2x2 + 34 is it one of theas anwsers

OpenStudy (stamp):

@sam612 The problem is solved. Do not ask for help and then ignore it when given...

OpenStudy (anonymous):

ok